OpenStudy (anonymous):

I am doing quadratic functions Can someone help me solve this y=-2x^2+11x-12

OpenStudy (anonymous):

Solve it for y = 0? Are you familiar with the quadratic formula?

OpenStudy (anonymous):

yes but this is +or - a=-2,b=11 c=-12

OpenStudy (anonymous):

I know how to do the formula but have done 10 problems but don't come up with the correct answer

OpenStudy (anonymous):

Those are the correct values for a, b, and c.

OpenStudy (anonymous):

\[11+-\sqrt{(11^2-4(-2)(-12)}\]

OpenStudy (anonymous):

I have done the numbers in parenthesis first what do you come up with

OpenStudy (anonymous):

\(11^2 - 4(-2)(-12) = 25\)

OpenStudy (anonymous):

okay that is what I got

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

So you should have \(\frac{-11 \pm \sqrt{25}}{2(-2)}\)

OpenStudy (anonymous):

I have to have a + and - answer

OpenStudy (anonymous):

They could both be + or - depending on your other values.

OpenStudy (anonymous):

But you will have 2 unless the part under the sqrt is 0

OpenStudy (anonymous):

I have \[11+-\sqrt{25?}\div-4\]

OpenStudy (anonymous):

It should be -11, not 11. Recall that the quadratic formula is \[ax^2 + bx +c=0 \implies x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

i got x=-9 and -14 over-4

OpenStudy (anonymous):

okay that is right I have it in front of me

OpenStudy (anonymous):

x = (-11 + 5 )/(-4) = -6/-4 = 3/2 x = (-11 - 5)/(-4) = -16/-4 = 4

OpenStudy (anonymous):

if you let y be 0 than after you can write your equation as -2x^2+11x-12=0 or you can write this as 2x^2-11x+12=0 from this equation you can get the correct answer

OpenStudy (anonymous):

You won't have any negative values for the x intercepts of this graph.

OpenStudy (anonymous):

True I have been working on this too long I think thank you

OpenStudy (anonymous):

Certainly =)